The following function defines a recursive sequence:
f(0) = -4
f(1) = 12
f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1
Which of the following sequences is defined by this recursive function?
-4, 12, -28, 60, …
-4, -12, -28, -60, …
-4, 12, -18, 54, …
-4, 12, -18, -54, …

Answer by jim_thompson5910(35256)


Plug in n = 2 to get
f(n) = -3*f(n -1) - 2*f(n - 2)
f(2) = -3*f(2 -1) - 2*f(2 - 2)
f(2) = -3*f(1) - 2*f(0)
f(2) = -3*12 - 2*(-4)
f(2) = -28

Plug in n = 3 to get
f(n) = -3*f(n -1) - 2*f(n - 2)
f(3) = -3*f(3 -1) - 2*f(3 - 2)
f(3) = -3*f(2) - 2*f(1)
f(3) = -3*(-28) - 2*(12)
f(3) = 60

So the sequence is -4, 12, -28, 60